[Hallc_running] Hall C Status

Howard Fenker hcf at jlab.org
Tue Dec 12 17:35:15 EST 2017

Hello everybody,

The magnet experts spent some hours in the Hall this afternoon to make 
the HMS Q1
controls more stable. We hope this eliminates some of the trips it has 
been suffering.

The target fan has issues (runs at 36 Hz). Target experts offered to try 
to fix this
but we declined so that we can continue collecting HMS data under 
consistent target

We should have beam back in the hall soon. It is scheduled to continue 
until 8:00 a.m.
tomorrow at which point work on the separators will interrupt beam to 
all halls. We do
anticipate getting it back during the day shift, however, so we need a 
full shift crew.

With the SHMS offline we are taking calibration data for the HMS. 
Getting the HMS back
into a known state will make it much easier to calibrate the SHMS when 
it rejoins us.


More information about the Hallc_running mailing list